# On the Hausdorff dimension and Cantor set structure of sliding Shilnikov invariant sets

### v2

@@ Abstract.
The concept of sliding Shilnikov connection has been recently introduced and represents an important notion in Filippov systems, because its existence implies chaotic behavior on an invariant subset of the system. The investigation of its properties has just begun, and understanding the topology and complexity of its invariant set is of interest. In this paper, we conduct a local analysis on the first return map associated to a Shilnikov sliding connection, which reveals a conformal iterated function system (CIFS) structure. By using the theory of CIFS, we estimate the Hausdorff dimension of the local invariant set of the first return map, showing, in particular, that it is strictly greater than 00 and strictly less than $1$, and its one-dimensional Lebesgue measure is 0. Moreover, we prove that the closure of the local invariant set is a Cantor set and has the same Hausdorff dimension and Lebesgue measure of the original invariant set. Furthermore, it is given by the invariant set adjoined with the set of all pre-images of the regular-fold point.
@@ 1. Introduction
When modeling various phenomena, it is often observed that the rules governing their evolution change abruptly at specific thresholds, introducing discontinuities in their models (see [10, 11] for a general discussion on discontinuities in applied models). These abrupt changes are commonly associated with processes involving decisions, switches like in neurons or electronic processes, light refraction, body collisions, changes in dry friction regimes, or any other scenarios that exhibit abrupt shifts in behavior. For example, in ecology, the phenomenon of prey-switching describes a predator’s adaptive diet in response to the availability of different prey species. This behavior, observed in many predator species, creates discontinuities in prey-predator models (see, for example, [24, 30]). Similar discontinuities are found in other applied models, such as prey-predator models with prey refuge [13], mechanical systems [28], electromagnetic processes [2], and others. The mathematical framework used to model and understand these phenomena includes the concept of piecewise smooth differential systems. Therefore, obtaining a better understanding of the geometric and dynamical properties of these systems is highly valuable. For surveys on piecewise smooth dynamical systems and their applications, see [3, 14].
@@
Filippov’s convention and its related concepts will be formally defined right below in the following subsections. But before that, let us briefly discuss, in an informal manner, this convention and the main object of our study. First, for points on the switching manifold where the vector fields in both sides cannot be concatenated to create a trajectory that crosses the switching manifold, the Filippov’s convention induces a dynamics on the switching set, allowing trajectories to slide along it, a phenomenon known as sliding dynamics. Also, the Filippov’s convention extends the classical concept of singularities for smooth vector fields by inducing new types of singularities on the switching set. In these scenarios, the trajectories of a Filippov system may asymptotically approach these singularities or reach them in finite time, sometimes leading to a loss of uniqueness of trajectories. The combination of these new singularities and the sliding dynamics gives rise to unique global phenomena in Filippov systems, such as the sliding Shilnikov connection (see Definition A) that has been recently introduced and discussed in [22].
Sliding Shilnikov connection is an important notion in Filippov systems, as their existence implies chaotic behavior within an invariant subset of the system, as demonstrated in [21]. This phenomenon has practical applications in applied science. For instance, in [24], numerical evidences of chaotic behavior were observed in a prey-switching Filippov-type prey-predator model. This observation was analytically explained in [4] by proving that such a model exhibits a sliding Shilnikov connection. The study of the properties and applicability of this connection is in its early stages, and understanding the topology and complexity of its associated invariant sets is of interest.
@@ 1.1. Filippov systems
where $X,Y\in\mathcal{C}^{l,\varepsilon}(V,\mathbb{R}^{n}),$ $l\geq 1$, and $g\in\mathcal{C}^{1}(V,\mathbb{R})$ is a function with 00 as a regular value (that is, $Dg(u):\mathbb{R}^{n}\to\mathbb{R}$ is surjective for all $u\in g^{-1}(0)$). Its switching manifold is given by $M=g^{-1}(0)$. Recall that a function $f:V\subseteq\mathbb{R}^{n}\to\mathbb{R}^{m}$ is said to be of class $\mathcal{C}^{l,\varepsilon}(V,\mathbb{R}^{m})$ if $f\in\mathcal{C}^{l}(V,\mathbb{R}^{m})$ and its $l$-th derivative, $D^{l}f$, satisfies the $\varepsilon$-Hölder condition. Specifically, there exists a constant $L$ such that $|D^{l}f(x)-D^{l}f(y)|\leq L|x-y|^{\varepsilon}$ for all $x,y\in V$. When the context is clear, we may abbreviate this as $\mathcal{C}^{l,\varepsilon}(V)$ or $\mathcal{C}^{l,\varepsilon}$.
The piecewise smooth vector field (1) is concisely denoted as $Z=(X,Y)_{g}$ (or simply $Z=(X,Y)$). The space of all piecewise smooth systems of the form (1) is denoted by $\Omega_{g}^{l,\varepsilon}(V,\mathbb{R}^{n})\cong\mathcal{C}^{l,\varepsilon}(V,\mathbb{R}^{n})\times\mathcal{C}^{l,\varepsilon}(V,\mathbb{R}^{n})$, allowing us to endow it with the product topology. When the context is clear, we may abbreviate this as $\Omega_{g}^{l,\varepsilon}(V)$ or $\Omega_{g}^{l,\varepsilon}$.
The Filippov’s convention establishes that the local trajectories of $Z$ (i.e. local solution of the differential system $\dot{u}=Z(u)$) correspond to solutions of the differential inclusion
$\dot{u}\in\mathfrak{F}(u),\,\,u\in V,$ (2)
where $\mathfrak{F}:V\leadsto\mathbb{R}^{n}$ is the following set-valued function
$\mathfrak{F}_{Z}(u):=\begin{cases}\left\{Y(u)\right\},&\quad\text{if }g(u)<0,\\ \left\{s\,X(u)+(1-s)Y(u):t\in[0,s]\right\},&\quad\text{if }g(u)=0,\\ \left\{X(u)\right\},&\quad\text{if }g(u)>0.\end{cases}$
We recall that $\varphi:I\to V$, defined on an open interval $I\subseteq\mathbb{R}$, is said to be a solution of the differential inclusion (2), if it is an absolutely continuous function satisfying $\dot{\varphi}(t)\in\mathfrak{F}(\varphi(t))$ for almost every $t\in I$. For an introduction on differential inclusions, see [1].
The local trajectories of (1) has an intuitive geometric interpretation. To explore this, let $\varphi^{t}_{F}(u)$ denote the flow of a vector field $F:V\subseteq\mathbb{R}^{n}\to\mathbb{R}^{n}$ at time $t$ starting from $u$, and also define $Fg(u):=\langle F(u),\nabla g(u)\rangle$. For points in $V$ where $g(u)\neq 0$, the local trajectory corresponds to the local trajectories of either $X$ or $Y$, depending on whether $g(u)>0$ or $g(u)<0$, respectively. To describe the local solutions for points on the switching manifold $M$, we first distinguish between some open regions on $M$ (see Figure 1):
@@
  • The points on $M$ satisfying $Xg(u)<0$ and $Yg(u)>0$ define the escaping region $M^{e}$. This implies that there exists $t_{2}>0$ such that $g(\varphi_{X}^{t}(u))>0$ for $t\in(0,t_{2})$ and $g(\varphi_{Y}^{t}(u))<0$ for $t\in(0,t_{2})$, so that both trajectories also cannot be concatenated at $u$ to form a trajectory of $Z$.
@@
Figure 1. A graphical representation of the crossing, sliding and escaping regions on the switching manifold.
@@
The sliding dynamics, described above, naturally introduces a first new type of singularity of the Filippov system $Z$, corresponding to the singularities of the sliding vector field $\tilde{Z}$. These are points $u^{*}\in M^{s,e}$ where $\tilde{Z}(u^{*})=0$, known as pseudo-equilibria of $Z$. Notably, trajectories of $\tilde{Z}$ can asymptotically approach a pseudo-equilibrium, while trajectories of $X$ and $Y$ can either reach or depart from it in finite time. This type of singularity is crucial in the definition of sliding Shilnikov orbits. A pseudo-equilibrium is considered hyperbolic if it is a hyperbolic singularity of $\tilde{Z}$. Additionally, if $u^{*}\in M^{s}$ is an unstable hyperbolic focus of $\tilde{Z}$, or if $u^{*}\in M^{e}$ is a stable hyperbolic focus of $\tilde{Z}$, then $u^{*}$ is referred to as a hyperbolic pseudo-saddle-focus (this corresponds to point $p$ in Figure 2).
Finally, we consider the set of tangency points $M^{t}$, which consists of points $u\in M$ where $Xg(u)Yg(u)=0$. A point in $M^{t}$ is called a tangency point of $X$ if $Xg(u)=0$, or a tangency point of $Y$ if $Yg(u)=0$. There are many possible configurations of points in $M^{t}$, leading to different definitions of their local trajectories. As a result, the uniqueness of local trajectories is also not guaranteed at points in $M^{t}$. In the following, we will introduce the concept of a visible fold-regular point, a specific type of tangency point that occurs in sliding Shilnikov connections.
A tangency point $u\in M$ is referred to as a visible fold of $X$ (resp. $Y$) if $X^{2}g(u):=X(Xg)(u)>0$ (resp. $Y^{2}g(u):=Y(Yg)(u)<0$). Conversely, if the inequalities are reversed, the point $u$ is called an invisible fold of $X$ (resp. $Y$). A visible/invisible fold $u\in M$ of $X$ (resp. $Y$) is called a visible/invisible fold-regular point if $Yh(u)\neq 0$ (resp. $Xh(u)\neq 0$). If $Yh(u)>0$ (resp. $Xh(u)<0$), the point lies on the boundary of the sliding region, $\partial M^{s}$. Conversely, if $Yh(u)<0$ (resp. $Xh(u)>0$), it lies on the boundary of the escaping region, $\partial M^{s}$. For instance, the point $q$ in Figure 2 corresponds to a visible fold-regular point lying on $\partial M^{s}$.
@@
    Firstly, the sliding vector field $\tilde{Z}$ is always transverse to these points, which means that the trajectories of $\tilde{Z}$ either reach or depart from them transversely within a finite time.
@@
    Secondly, when the dimension of the space is greater than or equal to $3$ ($n\geq 3$), a visible fold-regular point $q$ is never isolated, which means that there exists a neighborhood $U$ around $q$ such that $U\cap\partial M^{s,e}$ is a curve consisting entirely of visible fold-regular points.
@@ 1.2. Sliding Shilnikov connection
The concept of sliding Shilnikov connection was recently introduced in [22]. Roughly speaking, it consists of a trajectory $\Gamma$ of $Z$, passing though a visible fold-regular $q\in\partial M^{s}$ and connecting a hyperbolic pseudo-saddle point $p$ to itself asymptotically on one side and in finite time on the other side (see Figure 2).
Some of its dynamical properties, such as chaotic behavior, was further explored in [21]. It has also demonstrated significant applied importance, as shown in [4], by proving that a family of prey-switching Filippov-type prey-predator models exhibits chaotic behavior, which had previously been supported only by numerical evidence in [24].
@@
Let $Z=(X,Y)$ be a Filippov system with a hyperbolic pseudo saddle-focus $p\in M^{s}$ (resp. $p\in M^{e}$) and a visible fold-regular point $q\in\partial M^{s}$ (resp. $q\in\partial M^{e}$), which is a visible fold point of $X$. Assume that:
@@
    The trajectory of $\tilde{Z}$ passing through $q$ converges to $p$ backward in time (resp. forward in time), that is, $\lim_{t\to-\infty}\varphi^{t}_{\tilde{Z}}(q)=p$ (resp. $\lim_{t\to+\infty}\varphi^{t}_{\tilde{Z}}(q)=p$);
@@
Figure 2. Representation of a sliding Shilnikov connection. The point $p\in M^{s}$ is a hyperbolic pseudo saddle-focus and the point $q\in\partial M^{s}$ is a visible fold-regular point for $X$. The forward trajectory of $Z$ at $q$ follows the flow of the vector field $X$ until it reaches, in finite time, the sliding region $M^{s}$ at the hyperbolic pseudo saddle-focus $p$. The backward trajectory of $Z$ at $q$ follows the backward flow of the sliding vector field $\tilde{Z}$ which approaches asymptotically to the hyperbolic pseudo saddle-focus $q$.
In [21], it was demonstrated the existence of a neighborhood $B\subseteq\mathbb{R}^{3}$ of $q$ such that, for $\gamma:=B\cap\partial M^{s,e}$, which is a curve of visible fold-regular points (see Property (A2) from Remark 1), a first return map $\pi:\operatorname{Dom}(\pi)\to\gamma$ is well-defined on a subset $\text{Dom}(\pi)\subseteq\gamma$. This map captures the complete dynamics of the Filippov system $Z$ in a neighborhood of the sliding Shilnikov connection $\Gamma$. Focusing on the points in $\operatorname{Dom}(\pi)$, which are the ones that return to $\gamma$, the dynamics of $\pi$ exhibit a rich structure with many interesting properties. For instance, in [21], it was shown that the restriction of $\pi$ to an invariant set is topologically conjugate to a Bernoulli shift with infinite topological entropy. In Section 3.1, we will provide details on the construction of the first return map $\pi$.
@@
As we will show, the first return map can be decomposed into maps exhibiting expanding behavior, such that the functions corresponding to their inverses are contractions (see Figure 4). This brings us into the realm of Iterated Function Systems (IFS), or more specifically, Conformal Iterated Function Systems (CIFS) when the functions meet suitable criteria. We will use tools from IFS theory to analyze the geometric aspects of $\Lambda_{U}$ mentioned before (see Section 2.3)

### v4

@@ Abstract.
The concept of sliding Shilnikov connection has been recently introduced and represents an important notion in Filippov systems, because its existence implies chaotic behavior on an invariant subset of the system. The investigation of its properties has just begun, and understanding the topology and complexity of its invariant set is of interest. In this paper, we conduct a local analysis on the first return map associated to a sliding Shilnikov connection, which reveals a conformal iterated function system (CIFS) structure. By using the theory of CIFS, we estimate the Hausdorff dimension of the local invariant set of the first return map, showing, in particular, that is a positive number smaller than $1$, and with one-dimensional Lebesgue measure equal to 00. Moreover, we prove that the closure of the local invariant set is a Cantor set and retains both the Hausdorff dimension and Lebesgue measure of the local invariant set. Furthermore, this closure consists of the local invariant set along with the set of all pre-images, under the first return map, of the visible fold-regular point contained in the connection.
@@ 1. Introduction
When modeling various phenomena, it is often observed that the rules governing their evolution change abruptly at specific thresholds, introducing discontinuities in their models (see [10, 11] for a general discussion on discontinuities in applied models). These abrupt changes are commonly associated with processes involving decisions, or switches, such as those in neurons or electronic systems, light refraction, body collisions, changes in dry friction regimes, or any other scenarios that exhibit sudden shifts in behavior. In ecology, for example, the phenomenon of prey-switching describes a predator’s adaptive diet in response to the availability of different prey species. This behavior, observed in many predator species, creates discontinuities in prey-predator models (see, for example, [24, 30]). Similar discontinuities are found in other applied models, such as prey-predator models with prey refuge [13], mechanical systems [28], electromagnetic processes [2], and others. The mathematical framework used to model and understand these phenomena includes the concept of piecewise smooth differential systems. Therefore, obtaining a better understanding of the geometric and dynamical properties of these systems is highly valuable. For surveys on piecewise smooth dynamical systems and their applications, see [3, 14].
@@
Filippov’s convention and its related concepts will be formally defined in the following subsections but, before that, let us briefly discuss in an informal manner this convention and the main object of our study. First, for points on the switching manifold where the vector fields in both sides cannot be concatenated to create a trajectory that crosses the switching manifold, the Filippov’s convention induces a dynamics on the switching manifold, allowing trajectories to slide along it, a phenomenon known as sliding dynamics. Also, the Filippov’s convention extends the classical concept of singularities for smooth vector fields by inducing new types of singularities on the switching manifold. In these scenarios, the trajectories of a Filippov system may asymptotically approach these singularities or reach them in finite time, potentially leading to a loss of uniqueness of trajectories. The combination of these new singularities and the sliding dynamics gives rise to unique global phenomena in Filippov systems, such as the sliding Shilnikov connection (see Definition A) that has been recently introduced and discussed in [22].
The sliding Shilnikov connection is an important notion in Filippov systems, as their existence implies chaotic behavior within an invariant subset of the system, as demonstrated in [21]. This phenomenon has practical applications in applied science. For instance, in [24], numerical evidences of chaotic behavior were observed in a prey-switching Filippov-type prey-predator model. This observation was analytically explained in [4] by proving that such a model exhibits a sliding Shilnikov connection. The study of the properties and applicability of this connection is in its early stages, and understanding the topology and complexity of its associated invariant sets is of interest.
@@ 1.1. Filippov systems
[…]
The piecewise smooth vector field (1) is concisely denoted as $Z=(X,Y)_{g}$ (or simply $Z=(X,Y)$). The space of all piecewise smooth systems of the form (1) is denoted by $\Omega_{g}^{l,\varepsilon}(V,\mathbb{R}^{n})\cong\mathcal{C}^{l,\varepsilon}(V,\mathbb{R}^{n})\times\mathcal{C}^{l,\varepsilon}(V,\mathbb{R}^{n})$, allowing us to endow it with the product topology. When the context is clear, we may abbreviate this as $\Omega_{g}^{l,\varepsilon}(V)$, $\Omega_{g}^{l,\varepsilon},\Omega^{l,\varepsilon}(V)$, or simply $\Omega^{l,\varepsilon}$.
The Filippov’s convention establishes that the local trajectories of $Z$ (i.e. local solutions of the differential system $\dot{u}=Z(u)$) correspond to solutions of the differential inclusion
$\dot{u}\in\mathfrak{F}_{Z}(u),\,\,u\in V,$ (2)
where $\mathfrak{F}_{Z}:V\leadsto\mathbb{R}^{n}$ is the following set-valued function
$\mathfrak{F}_{Z}(u):=\begin{cases}\left\{X(u)\right\},&\quad\text{if }g(u)>0,\\ \left\{(1-s)X(u)+s\,Y(u):t\in[0,s]\right\},&\quad\text{if }g(u)=0,\\ \left\{Y(u)\right\},&\quad\text{if }g(u)<0.\end{cases}$
We recall that $\varphi:I\to V$, defined on an open interval $I\subseteq\mathbb{R}$, is said to be a solution of the differential inclusion (2), if it is an absolutely continuous function satisfying $\dot{\varphi}(t)\in\mathfrak{F}_{Z}(\varphi(t))$ for almost every $t\in I$. For an introduction on differential inclusions, see [1].
The local trajectories of (1) have an intuitive geometric interpretation. To explore this, let $\varphi^{t}_{F}(u)$ denote the flow of a vector field $F:V\subseteq\mathbb{R}^{n}\to\mathbb{R}^{n}$ at time $t$ starting from $u$, and also define
$Fg(u):=\langle F(u),\nabla g(u)\rangle,$ (3)
[…]
@@
  • The points on $M$ satisfying $Xg(u)>0$ and $Yg(u)<0$ define the escaping region $M^{e}$. This implies that there exists $t_{2}>0$ such that $g(\varphi_{X}^{t}(u))>0$ for $t\in(0,t_{2})$ and $g(\varphi_{Y}^{t}(u))<0$, so that both trajectories also cannot be concatenated at $u$ to form a trajectory of $Z$.
@@
Figure 1. A graphical representation of the crossing, escaping and sliding regions on the switching manifold.
@@
The sliding dynamics, described above, naturally introduces a first new type of singularity of the Filippov system $Z$, corresponding to the singularities of the sliding vector field $\tilde{Z}$. These are points $u^{*}\in M^{s,e}$ where $\tilde{Z}(u^{*})=0$, known as pseudo-equilibria of $Z$. Notably, trajectories of $\tilde{Z}$ can asymptotically approach a pseudo-equilibrium, while trajectories of $X$ and $Y$ can either reach or depart from it in finite time. This type of singularity is crucial to the definition of sliding Shilnikov connections. A pseudo-equilibrium is considered hyperbolic if it is a hyperbolic singularity of $\tilde{Z}$. Additionally, if $u^{*}\in M^{s}$ is an unstable hyperbolic focus of $\tilde{Z}$, or if $u^{*}\in M^{e}$ is a stable hyperbolic focus of $\tilde{Z}$, then $u^{*}$ is referred to as a hyperbolic pseudo-saddle-focus (this corresponds to the point $p$ in Figure 2).
Finally, we consider the set of tangency points $M^{t}$, which consists of points $u\in M$ where $Xg(u)Yg(u)=0$. A point $u\in M^{t}$ is called a tangency point of $X$ if $Xg(u)=0$, or a tangency point of $Y$ if $Yg(u)=0$. There are many possible configurations of points in $M^{t}$, leading to different definitions of their local trajectories. As a result, the uniqueness of local trajectories is also not guaranteed at points in $M^{t}$. In the following, we will introduce the concept of a visible fold-regular point, a specific type of tangency point that occurs in sliding Shilnikov connections.
A tangency point $u\in M^{t}$ is referred to as a visible fold of $X$ (resp. $Y$) if $X^{2}g(u):=X(Xg)(u)>0$ (resp. $Y^{2}g(u):=Y(Yg)(u)<0$). Conversely, if the inequalities are reversed, the point $u$ is called an invisible fold of $X$ (resp. $Y$). A visible/invisible fold $u\in M$ of $X$ (resp. $Y$) is called a visible/invisible fold-regular point if $Yg(u)\neq 0$ (resp. $Xg(u)\neq 0$). If $Yg(u)>0$ (resp. $Xg(u)<0$), the point lies on the boundary of the sliding region, $\partial M^{s}$. Conversely, if $Yg(u)<0$ (resp. $Xg(u)>0$), it lies on the boundary of the escaping region, $\partial M^{e}$. For instance, the point $q$ in Figure 2 corresponds to a visible fold-regular point lying on $\partial M^{s}$.
@@
    Firstly, the sliding vector field $\tilde{Z}$ is always transverse to these points, which means that trajectories of $\tilde{Z}$ either reach or depart from them transversely within a finite time.
@@
    Secondly, when the dimension of the space is greater than or equal to $3$ ($n\geq 3$), a visible fold-regular point $q$ is never isolated, which means that there exists a neighborhood $U$ around $q$ such that $U\cap\partial M^{s,e}$ is a set consisting entirely of visible fold-regular points.
@@ 1.2. Sliding Shilnikov connection
The concept of sliding Shilnikov connection was recently introduced in [22]. Roughly speaking, it consists of a trajectory $\Gamma$ of $Z$, passing though a visible fold-regular point $q\in\partial M^{s,e}$ and connecting a hyperbolic pseudo-saddle-focus $p$ to itself asymptotically on one side and in finite time on the other side (see Figure 2).
Some of its dynamical properties, such as chaotic behavior, were further explored in [21]. It has also demonstrated significant applied importance, as shown in [4], by proving that a family of prey-switching Filippov-type prey-predator models exhibits chaotic behavior, which had previously been supported only by numerical evidence in [24].
@@
Let $Z=(X,Y)\in\Omega^{1,\varepsilon}$ be a Filippov system with a hyperbolic pseudo-saddle-focus $p\in M^{s}$ (resp. $p\in M^{e}$) and a visible fold-regular point $q\in\partial M^{s}$ (resp. $q\in\partial M^{e}$), which is a visible fold point of $X$. Assume that:
@@
    The trajectory of $\tilde{Z}$ passing through $q$ converges to $p$ backward in time (resp. forward in time), that is, $\lim_{t\to-\infty}\varphi^{t}_{\tilde{Z}}(q)=p$ (resp. $\lim_{t\to+\infty}\varphi^{t}_{\tilde{Z}}(q)=p$).
@@
Figure 2. Representation of a sliding Shilnikov connection. The point $p\in M^{s}$ is a hyperbolic pseudo-saddle-focus and the point $q\in\partial M^{s}$ is a visible fold-regular point for $X$. The forward trajectory of $Z$ at $q$ follows the flow of the vector field $X$ until it reaches, in finite time, the sliding region $M^{s}$ at the hyperbolic pseudo-saddle-focus $p$. The backward trajectory of $Z$ at $q$ follows the backward flow of the sliding vector field $\tilde{Z}$ which approaches asymptotically to the hyperbolic pseudo-saddle-focus $p$.
In [21], it was demonstrated the existence of a neighborhood $B\subseteq\mathbb{R}^{3}$ of $q$ such that, for $\gamma:=B\cap\partial M^{s,e}$, which is a curve of visible fold-regular points (see (A2) from Remark 1), a first return map $\pi:\operatorname{Dom}(\pi)\to\gamma$ is well-defined on a subset $\text{Dom}(\pi)\subseteq\gamma$. This map captures the complete dynamics of the Filippov system $Z$ in a neighborhood of the sliding Shilnikov connection $\Gamma$. The dynamics of $\pi$ exhibits a rich structure with many interesting properties. For instance, in [21], it was shown that the restriction of $\pi$ to an invariant set is topologically conjugate to a Bernoulli shift with infinite topological entropy. In Section 3.1, we will provide details on the construction of the first return map $\pi$.
@@
As we will show, the first return map can be decomposed into maps exhibiting expanding behavior, such that the functions corresponding to their inverses are contractions (see Figure 4). This brings us into the realm of Iterated Function Systems (IFS), or more specifically, Conformal Iterated Function Systems (CIFS) when the functions meet suitable criteria. We will use tools from IFS theory to analyze the geometric aspects of $\Lambda_{U}$ mentioned before (see Section 2.3).